A weight is suspended from a spring and is moving up and down in a simple harmonic motion. At start, the weight is pulled down 5 cm below the resting position, and then released. After 8 seconds, the weight reaches its highest location for the first time. Find the equation of the motion. 


1
Expert's answer
2022-01-30T16:05:31-0500


General equation for simple Harmonic equation of spring is;


"X=" A cos "(2\pi\>ft+\phi)"

"X=" Displacement

"A=" Amplitude "=0.05m"

"f=" Frequency

"t=" time


and "\phi=" Phase angle


at "t=0,x=-5"


"\therefore-0.05=0.05\>cos(0+\phi)"

"\implies\phi=\pi"



Period time T "=8×2=16\>" seconds

"f=\frac{1}{T}=0.0625"


"\therefore\>X=0.05\>cos(0.3927t+\pi)"